A sociologist develops a tost to measure ormally dstbuted. Find the 96% confidence nterv㎡ for the mean score of all such subjects Rond toone decimal place O A 092 to 83.2 O B. 74.6 to 77.8 C. 642to882 O D. 67.7 to 84.7
Answer #1

SE=s/sqrt{n}

= 21.4/V27 = 4.1184

alpha = 0.05

ndf = 27 - 1 = 26

From Table, critical values of t = pm 2.0556

Confidence interval:

76.2 pm (2.0556 X 4.1184)

= 76.2 pm 8.4658

= ( 67.7, 84.7)

So,

correct option:

D. 67.7 to 84.7